lazyking4life:

Given m \| nm∥n, find the value of x. (6x+3)=(2x+13)

supie:

Ok so the first step would be to subtract `2x` from both sides of the equation do you think you can do that?

lazyking4life:

no bc what side n what numbers

supie:

I'm confused- I dont think I've learned this @jhonyy9 any idea ?

Florisalreadytaken:

hey @lazyking4life can we have a screensnip of that?

lazyking4life:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Florisalreadytaken hey @lazyking4life can we have a screensnip of that? \(\color{#0cbb34}{\text{End of Quote}}\) Ok

lazyking4life:

Florisalreadytaken:

ah ok that's better -- imma help you in a bit ok

lazyking4life:

Ok

Florisalreadytaken:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @lazyking4life \(\color{#0cbb34}{\text{End of Quote}}\) ok, so I'm seeing that you've inserted 2.5 as an answer -- why is that?

lazyking4life:

I don’t mean to type that

Florisalreadytaken:

(6x+3)=(2x+13) tis is wrong -- its (6x-3)=(2x+13)

Florisalreadytaken:

this*

Florisalreadytaken:

firstly we move whatever contains x to the left side

Florisalreadytaken:

ehm just do the math its that easy

Florisalreadytaken:

tell me if you need help with that

KynxKatz:

Do u still need help frenn?

lazyking4life:

Yes but nomo

jhonyy9:

opposite angles are congruent use this - so what can you write ?

jhonyy9:

6x-3=2x+13 solve this for x

jhonyy9:

like a first step subtract from both sides 2x so what will get ?

jhonyy9:

lazyking dont wan collaborate @supie do you wan finish it ?

jhonyy9:

are you lazyking ?